OpenStudy (anonymous):

A LITTLE CONFUSED Find the y value for the point that splits segment EF in half if point E is located at (5, -2) and point F is located at (-1.3)

OpenStudy (anonymous):

Find the mid point between E and F

OpenStudy (anonymous):

3- (-2) ? is that how you find it ? @nelsonjedi

OpenStudy (anonymous):

Mid point x= (x1+x2) / 2 y= (y1+y2) / 2

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

You are welcome